• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name V. Sulkosky

    Log entry time 18:57:53 on September 14, 2008

    Entry number 243182

    keyword=Turned on LHRS VDC Discriminator Thresholds

    We traced the deadtime problem in the LHRS to ROC 3. The environment
    became noisy and the TDCs were seeing numerous hits causing the DAQ to
    slow down considerably. Bob turned on the Left VDC discriminator
    LVL to 4.0 V. However this did not fix the problem. I noticed that the
    voltage levels were 0. V, when the Tools GUI indicates they should
    be -5.2 V and +5 V. I conferred with Jack and he said that the power
    supplies that drive the thresholds were still off.

    Once we turned the power supplies on, I checked the DAQ with a T8 rate of
    500 Hz and saw no deadtime for over 10 mintues. In the past, the deadtime
    would climb up to 98% within about 5 minutes of running.

    The DAQ appears to be working now at least until the next problem!

    A copy of this log entry has been emailed to: rom@jlab.org,segal@jlab.org